# and Machine Learning

### v3

@@ 3.3 Kernel Denoisers
The basic idea behind kernel denoisers follows a non-parametric approach to modeling the distribution of (clean) images. Concretely, consider our basic setting
$\mathbf{x}=\mathbf{u}+\mathbf{e},$ (31)
@@
$\widehat{P}(\mathbf{u})=\frac{1}{n}\sum_{i=1}^{n}\delta(\mathbf{u}-\mathbf{u}_{i})$
@@
$\widehat{P}(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})$
@@
$\nabla\log\widehat{P}(\mathbf{x})=\frac{-1}{\sigma^{2}}\left[\mathbf{x}-\frac{\sum_{i}\mathbf{u}_{i}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})}{\sum_{i}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})}\right].$ (32)
Invoking Tweedie’s formula, we have a closed form approximation to the MMSE denoiser as a (data-dependent) weighted average of the clean data points $\mathbf{u}_{i}$:
@@
$=\sum_{i}\mathbf{u}_{i}W(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})$
@@
$\hat{x}_{i}=\sum_{i}x_{j}W(x_{j}-x_{i},\alpha).$
This is a primitive instance of the pseudo-linear form alluded to earlier. In particular, the Gaussian ”kernels”, motivated by the assumed (Gaussian) distribution of the noise, can be thought of more generally as one of a myriad of choices of positive-definite kernels that can be employed to construct more general denoisers, as described below.
@@
Generally speaking, the weights are computed based on the affinity (or similarity) of pixels, measured using a ”kernel” (a symmetric positive-definite function). When properly normalized, these yield the weights used to compute the output pixels as a weighted average. For instance, in the non-local means [35] case
$K_{ij}(\mathbf{x},\alpha)=\exp(-\|\mathbf{R}_{ij}\mathbf{x}\|^{2}/2\alpha^{2}),\;\;\;\;\;\text{where}\;\;\;\;\|\mathbf{R}_{ij}\mathbf{x}\|^{2}=\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}$ (36)
@@
$\mathbf{W}(\mathbf{x},\alpha)=\mathbf{D}^{-1}(\mathbf{x},\alpha)\mathbf{K}(\mathbf{x},\alpha)$ (38)
@@
$f(\mathbf{x},\alpha)=\nabla\left[\mathbf{x}^{T}\mathbf{W}\mathbf{x}\right]$ (39)
In practice, symmetry of the filter matrix $\mathbf{W}(\mathbf{x},\alpha)$ is not a given [^12] Though one can empirically verify that such weight matrices are approximately symmetric [37, 38]. Despite the fact that the kernel matrix $\mathbf{K}(\mathbf{x},\alpha)$ is symmetric, the resulting weight matrix $\mathbf{W}(\mathbf{x},\alpha)=\mathbf{D}^{-1}(\mathbf{x},\alpha)\mathbf{K}(\mathbf{x},\alpha)$ is not so, due to the non-trivial diagonal normalization by $\mathbf{D}$. Fortunately, one can modify $\mathbf{W}(\mathbf{x},\alpha)$ to satisfying the symmetry condition as detailed in [1, 39]. This is accompliched by applying Sinkhorn balancing to $\mathbf{W}$ (or equivalently to $\mathbf{K}$), resulting in a symmetric and doubly-stochastic weight matrix, which can incidentally improve mean-squared error denoising performance over the baseline.
@@ 3.4 Summary
Shrinkage Estimators: The general form $f(\mathbf{x})=\mathbf{x}-\alpha v(\mathbf{x})$ can be interpreted as the ”trivial” denoiser $\mathbf{x}$ with a correction term $\alpha v(\mathbf{x})$ that pulls the components of the noisy input toward zero. It is remarkable that these denoisers have the same form as the original James-Stein estimator [41], where $\mathbf{x}$ was interpreted as the maximum-likelihood estimator, and $\alpha v(\mathbf{x})$ played the role of a Bayesian ”correction”. It has been observed [7, 30, 31] that such denoisers behave (at least locally) as shrinkage operations in an orthonormal basis adapted to the underlying structure of the image.
Gradient Descent on Energy: We noted that many denoisers can be written in the form $f(\mathbf{x})=\mathbf{x}-\alpha\nabla\mathcal{E}(\mathbf{x})$. It is obvious that the right-hand side defines one step in a steepest descent iteration. Repeated applications of a denoiser have the effect of marching toward the local stationary point of the energy.
Approximate Projection: It has been pointed out elsewhere [42] that if we accept the assertion that real-world images with $N$ pixels are approximately contained in low-dimensional manifolds of $\mathcal{R}^{N}$ [43], then adding noise is equivalent to orthogonal perturbation away from the manifold, and denoising is approximately a projection onto the manifold. In particular, for small noise, denoising is precisely a projection onto the local tangent of said manifold. As such, the work of denoising is essentially analogous to manifold learning.
@@ 4 Denoising, the Score Function, and Generative Modeling
A crucial link between denoising and the score function enables denoisers to enable learning complex probability distributions. In modeling real-world data, and images in particular, we are typically faced with a complex, high-dimensional probability density, $P(\cdot)$. Explicitly modeling such a distributions can be computationally intractable or extremely difficult. The score function, defined as the gradient of the log probability density, can provide a way through.
$s(\mathbf{x},\alpha)=\nabla_{\mathbf{x}}\log P(\mathbf{x},\alpha).$ (41)
@@
$s(\mathbf{x},\alpha)\approx\frac{\mathbf{x}-f(\mathbf{x},\alpha)}{\alpha}.$ (42)
At first blush, it is not at all clear why this is a reasonable procedure. Yet there are a number of ways [45, 46, 44, 47, 48] to motivate this idea -perhaps none more direct than by using Tweedie’s formula introduced earlier in Eq. (19).
@@
$f(\mathbf{x},\alpha)=\nabla\mathcal{E}_{0}(\mathbf{x},\alpha).$
@@
This illustrates again that the energy function is a proxy [49] for the score $s(\mathbf{x},\alpha)\approx\nabla\mathcal{E}(\mathbf{x},\alpha)$, and the resulting denoiser’s residual can be used as an approximation of the score.
@@ 4.1 Denoising as the Engine of Diffusion Models
Denoising Diffusion Models (DDMs) [46, 44, 51, 52, 48] have become an important area of research in generative modeling. They operate by progressively corrupting training data with noise until it’s indistinguishable from random noise, then learning to systematically reverse this corruption. By training a model to iteratively denoise, it gains the ability to generate entirely new, coherent data samples from a starting point of pure noise, effectively converting noise into meaningful structures like images or other data forms (Figure 6).
@@
$\mathbf{x}_{t}-\mathbf{x}_{t-dt}=\frac{\alpha_{t}-\alpha_{t-dt}}{2\alpha_{t}}\left(\mathbf{x}_{t}-\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]\right).$ (50)
From this, we can derive:
$\text{Var}[\mathbf{x}_{t-dt}|\mathbf{x}_{0}]$ $=\frac{(\alpha_{t}+\alpha_{t-dt})^{2}}{4\alpha_{t}^{2}}\text{Var}[\mathbf{x}_{t}|\mathbf{x}_{0}]$ (51)
$=\frac{(\alpha_{t}+\alpha_{t-dt})^{2}}{4\alpha_{t}}$ (52)
@@
$\approx\alpha_{t-dt}.$ (54)
@@ 5 Denoisers in the Context of Inverse Problems
Consider the following formulation of a linear inverse problem: The data is given by the following model $\mathbf{y}=H\mathbf{x}+\mathbf{e}$. A nominal solution can be obtained by solving this optimization problem:
$\hat{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{x}}l(\mathbf{y},\mathbf{x})+\lambda\mathbf{R}(\mathbf{x},\alpha)$ (55)
@@
More recently, and independently of the machine learning literature, a fascinating connection has emerged between denoising algorithms and inverse problems. Powerful denoising algorithms, particularly those leveraging deep learning, have been shown to implicitly encode strong priors about natural signals. By incorporating these denoisers into the optimization framework, we can effectively leverage their learned priors to achieve state-of-the-art performance in various inverse problems  [67, 68, 69, 70, 71, 37, 72, 73]. This approach effectively blurs the lines between traditional regularization techniques and modern denoising methods, offering a new paradigm for solving inverse problems.
@@
$H^{T}(\mathbf{y}-H\mathbf{x})+\lambda\nabla\mathbf{R}(\mathbf{x},\alpha)$ (56)
A key concern is how to compute $\nabla\mathbf{R}(\mathbf{x},\alpha)$. In this respect, classical choices of $\mathbf{R}(.)$ such as $L_{p}$ norms have been fairly convenient and successful; but also shown to have limited power to model natural images.
@@
$\mathbf{R}(\mathbf{x},\alpha)=\frac{1}{2}\mathbf{x}^{T}\;\mathbf{L}(\mathbf{x},\alpha)\;\mathbf{x}$ (57)
In [37], we developed a natural extension of this idea called Regularization by Denoising (RED), where the regularizer is constructed from a more general denoiser:
@@
$\nabla\mathbf{x}^{T}(\mathbf{x}-f(\mathbf{x},\alpha))$ $=$ $2\mathbf{x}-\nabla\left[\mathbf{x}^{T}f(\mathbf{x},\alpha)\right]$
$=$ $2\mathbf{x}-f(\mathbf{x},\alpha)-\nabla f(\mathbf{x},\alpha)\mathbf{x}$
$=$ $2\mathbf{x}-2f(\mathbf{x},\alpha)$
@@
$=$ $\lim_{\epsilon\rightarrow 0}\frac{(1+\epsilon)f(\mathbf{x})-f(\mathbf{x})}{\epsilon}=f(\mathbf{x})$ (60)
Replacing
$\nabla\mathbf{R}_{red}(\mathbf{x},\alpha)=\mathbf{x}-f(\mathbf{x},\alpha)$
for the gradient in (56), we have the following expression for the gradient of the objective:
$H^{T}(\mathbf{y}-H\mathbf{x})+\lambda(\mathbf{x}-f(\mathbf{x},\alpha))$
@@
$\mathbf{x}_{k+1}$ $=\left[H^{T}H+\lambda I\right]^{-1}H^{T}\mathbf{y}+\lambda\left[H^{T}H+\lambda I\right]^{-1}f(\mathbf{x}_{k},\alpha)$ (62)
$\mathbf{x}_{k+1}$ $=\mathbf{b}+\mathbf{M}\;f(\mathbf{x}_{k},\alpha)$ (63)
where $\mathbf{b}$ is the (fixed) linear pseudo-inverse solution and $\mathbf{M}$ is also a fixed matrix. Procedurally, we start with $\mathbf{x}_{0}=\mathbf{y}$, denoise it, and then a linear operator $\mathbf{M}$ is applied and a bias is added - this leads to an updated estimate, and the process is repeated. Note that the structure of this iterative process is not altogether different from a denoising diffusion process[51] where a denoiser is repeatedly applied. In fact, when $H=I$ we see the structure of a bridge diffusion process [78]:
$\mathbf{x}_{k+1}$ $=\frac{1}{1+\lambda}\mathbf{y}+\frac{\lambda}{1+\lambda}f(\mathbf{x}_{k},\alpha)$ (64)
@@
$\mathbf{R}(\mathbf{x},\alpha)=-\log P(\mathbf{x},\alpha),$ (65)
@@
$P(\mathbf{x},\alpha)\propto\exp{\left[-\mathbf{R}(\mathbf{x},\alpha)\right]},$ (66)
@@
An alternative approach to solving inverse problems involves generating samples from the posterior distribution using a pretrained prior in the form of a denoiser [80, 81, 82]. In other words, given measurements $\mathbf{y}$, we aim to generate samples $\mathbf{x}$ that follow the distribution $P(\mathbf{x}|\mathbf{y})$. We can achieve posterior sampling from $P(\mathbf{x}|\mathbf{y})$ by modifying equation (48) as follows:
$\frac{d\mathbf{x}_{t}}{dt}=-\frac{1}{2}\frac{d\alpha_{t}}{dt}\nabla\log P(\mathbf{x}_{t}|\mathbf{y},\alpha_{t})=-\frac{1}{2}\frac{d\alpha_{t}}{dt}\left(\nabla\log P(\mathbf{x}_{t},\alpha_{t})+\nabla\log P(\mathbf{y}|\mathbf{x}_{t},\alpha_{t})\right),$ (68)
starting from $x_{T}\sim P(x_{T})$. The second equality is given by Bayes rule. We recognized the score function $\nabla\log P(\mathbf{x}_{t},\alpha_{t})$, that can be again related to the MMSE denoiser using Tweedie’s formula (19). The second term captures how much the current $\mathbf{x}_{t}$ explains the measurements $\mathbf{y}$ and it is generally intractable. As an example, let us briefly present DPS [82]. In [82], authors propose to approximate the intractable term with $\log P(\mathbf{y}|\mathbf{x}_{t})\simeq\log P(\mathbf{y}|\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}])$, which is a consequence of assuming $p(\mathbf{x}_{0}|\mathbf{x}_{t})\simeq\delta(\mathbf{x}_{0}-\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}])$. This leads to a tractable formulation:
$\frac{d\mathbf{x}_{t}}{dt}=-\frac{1}{2}\frac{d\alpha_{t}}{dt}\left(\frac{\mathbf{x}_{t}-\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]}{\alpha_{t}}+\rho_{t}\nabla\|H\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]-\mathbf{y}||^{2}\right),$ (69)
where $\rho_{t}$ is a hyperparameter controlling the balance between the prior and the measurements.
Denoising diffusion models are rapidly gaining attention as a powerful tool for solving inverse problems in various fields. This is often achieved by combining the strengths of diffusion models with additional approximations or techniques, as evidenced by a growing body of research (e.g., [83, 48, 84, 80, 81, 85, 86, 87, 88, 89, 90]).

### v4

@@ 3.3 Kernel Denoisers
The basic idea behind kernel denoisers follows a non-parametric approach to modeling the distribution of (clean) images. Concretely, consider our basic setting given by
$\mathbf{x}=\mathbf{u}+\mathbf{e},$
@@
$\widehat{P}(\mathbf{u})=\frac{1}{n}\sum_{i=1}^{n}\delta(\mathbf{u}-\mathbf{u}_{i}).$ (34)
@@
$\widehat{P}(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I}).$ (35)
@@
$\nabla\log\widehat{P}(\mathbf{x})=-\frac{1}{\sigma^{2}}\left[\mathbf{x}-\frac{\sum_{i}\mathbf{u}_{i}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})}{\sum_{i}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})}\right].$ (36)
Invoking Tweedie’s formula, we have a closed form approximation to the MMSE denoiser as a (data-dependent) weighted average [3] of the clean data points $\mathbf{u}_{i}$:
@@
$=\sum_{i}\mathbf{u}_{i}W(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I}).$
@@
$\hat{x}_{i}=\sum_{i}x_{j}W(x_{j}-x_{i},\alpha).$ (38)
This is a primitive instance of the pseudo-linear form alluded to earlier. In particular, the Gaussian “kernels”, motivated by the assumed (Gaussian) distribution of the noise, can be thought of more generally as one of a myriad of choices of positive-definite kernels that can be employed to construct more general denoisers, as described below.
@@
Generally speaking, the weights are computed based on the affinity (or similarity) of pixels, measured using a “kernel” (a symmetric positive-definite function). When properly normalized, these yield the weights used to compute the output pixels as a weighted average. For instance, in the non-local means [59] case
$K_{ij}(\mathbf{x},\alpha)=\exp(-\|\mathbf{R}_{ij}\mathbf{x}\|^{2}/2\alpha^{2}),\;\;\;\;\;\text{where}\;\;\;\;\|\mathbf{R}_{ij}\mathbf{x}\|^{2}=\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2},$ (41)
@@
$\mathbf{W}(\mathbf{x},\alpha)=\mathbf{D}^{-1}(\mathbf{x},\alpha)\mathbf{K}(\mathbf{x},\alpha),$ (43)
@@
$f(\mathbf{x},\alpha)=\nabla\left[\mathbf{x}^{T}\mathbf{W}\mathbf{x}\right].$ (44)
In practice, symmetry of the filter matrix $\mathbf{W}(\mathbf{x},\alpha)$ is not a given [^14] Though one can empirically verify that such weight matrices are approximately symmetric [61, 62]. Despite the fact that the kernel matrix $\mathbf{K}(\mathbf{x},\alpha)$ is symmetric, the resulting weight matrix $\mathbf{W}(\mathbf{x},\alpha)=\mathbf{D}^{-1}(\mathbf{x},\alpha)\mathbf{K}(\mathbf{x},\alpha)$ is not so, due to the non-trivial diagonal normalization by $\mathbf{D}$. Fortunately, one can modify $\mathbf{W}(\mathbf{x},\alpha)$ to satisfying the symmetry condition as detailed in  [1, 63]. This is accomplished by applying Sinkhorn balancing to $\mathbf{W}$ (or equivalently to $\mathbf{K}$), resulting in a symmetric and doubly-stochastic weight matrix, which can incidentally improve mean-squared error denoising performance over the baseline - see also [64].
@@
To give some additional context to this approach, note that when applying a filter to an image, standard practice is to normalize the filter coefficients in order to maintain the local brightness level from input to output image. This is particularly important where nonlinear filters are concerned, where the effect on local brightness and contrast can be complex. The symmetrization approach presents a way of achieving the same level of control over the local filter behavior without the need for this normalization.
As described in [17], the approximation works better - in terms of the distortion introduced to the output image - when the diagonal entries of the matrix $\mathbf{D}$ are more tightly concentrated around their mean.
@@ 3.4 Summary
Shrinkage Estimators: The general form $f(\mathbf{x})=\mathbf{x}-\alpha v(\mathbf{x})$ can be interpreted as the “trivial” denoiser $\mathbf{x}$ with a correction term $\alpha v(\mathbf{x})$ that pulls the components of the noisy input toward zero. It is remarkable that these denoisers have the same form as the original James-Stein estimator [66], where $\mathbf{x}$ was interpreted as the maximum-likelihood estimator, and $\alpha v(\mathbf{x})$ played the role of a Bayesian “correction”. It has been observed [7, 54, 55] that such denoisers behave (at least locally) as shrinkage operations in an orthonormal basis adapted to the underlying structure of the image.
Gradient Descent on Energy: We noted that many denoisers can be written in the form $f(\mathbf{x})=\mathbf{x}-\alpha\nabla\mathcal{E}(\mathbf{x})$. It is obvious that the right-hand side defines one step in a steepest descent iteration. Repeated applications of a denoiser have the effect of marching toward a local stationary point of the energy.
Approximate Projection: It has been pointed out elsewhere [67] that if we accept the assertion that real-world images with $N$ pixels are approximately contained in low-dimensional manifolds of $\mathbb{R}^{N}$ [68], then adding noise is equivalent to orthogonal perturbation away from the manifold, and denoising is approximately a projection onto the manifold. In particular, for small noise, denoising is precisely a projection onto the local tangent of said manifold. As such, the work of denoising is essentially analogous to manifold learning.
@@ 4 Denoising, the Score Function, and Generative Modeling
A crucial link between denoising and the score function enables denoisers to learn complex probability distributions. In modeling real-world data, and images in particular, we are typically faced with a complex, high-dimensional probability density, $P(\cdot)$. Explicitly modeling such a distribution can be computationally intractable or extremely difficult. The score function, defined as the gradient of the log probability density, can provide a way through.
$s(\mathbf{x},\alpha)=\nabla\log P(\mathbf{x},\alpha).$ (46)
@@
$-s(\mathbf{x},\alpha)\approx\frac{\mathbf{x}-f(\mathbf{x},\alpha)}{\alpha}.$ (47)
At first blush, it is not at all clear why this is a reasonable procedure. Yet there are a number of ways [70, 71, 69, 72, 73] to motivate this idea -perhaps none more direct than by using Tweedie’s formula introduced earlier in Eq. (21):
@@
$f(\mathbf{x},\alpha)=\nabla\mathcal{E}_{0}(\mathbf{x},\alpha),$ (50)
@@
This illustrates again that the energy function is a proxy [74] for the score $s(\mathbf{x},\alpha)\approx-\nabla\mathcal{E}(\mathbf{x},\alpha)$, and the resulting denoiser’s residual can be used as an approximation of the score.
@@ 4.1 Denoising as the Engine of Diffusion Models
Denoising Diffusion and Flow generative models [71, 69, 76, 77, 73, 78] have become an important area of research in generative modeling. They operate by progressively corrupting training data with noise until it’s indistinguishable from random noise, then learning to systematically reverse this corruption. By training a model to iteratively denoise, it gains the ability to generate entirely new, coherent data samples from a starting point of pure noise, effectively converting noise into meaningful structures like images or other data forms (Figure 6).
@@
$\mathbf{x}_{t}-\mathbf{x}_{t-dt}$ $=\frac{\alpha_{t}-\alpha_{t-dt}}{2\alpha_{t}}\left(\mathbf{x}_{t}-\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]\right)$ (56)
$\mathbf{x}_{t-dt}$ $=\frac{\alpha_{t}+\alpha_{t-dt}}{2\alpha_{t}}\mathbf{x}_{t}-\frac{\alpha_{t}-\alpha_{t-dt}}{2\alpha_{t}}\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}].$ (57)
This allows us to derive the conditional variance of $\mathbf{x}_{t-dt}$ given $\mathbf{x}_{0}$:
$\text{Var}[\mathbf{x}_{t-dt}|\mathbf{x}_{0}]$ $=\text{Var}\left[\frac{\alpha_{t}+\alpha_{t-dt}}{2\alpha_{t}}\mathbf{x}_{t}-\frac{\alpha_{t}-\alpha_{t-dt}}{2\alpha_{t}}\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]\bigg{|}\mathbf{x}_{0}\right]$ (58)
$=\frac{(\alpha_{t}+\alpha_{t-dt})^{2}}{4\alpha_{t}^{2}}\text{Var}[\mathbf{x}_{t}|\mathbf{x}_{0}]$ (59)
$=\frac{(\alpha_{t}+\alpha_{t-dt})^{2}}{4\alpha_{t}^{2}}\cdot\alpha_{t}$ (60)
@@
$\approx\alpha_{t-dt},$ (62)
where the final approximation holds for small $dt$. This demonstrates that the velocity coefficient in (55) effectively ensures the consistent evolution of the conditional variance, crucial for accurately capturing the underlying process dynamics.
@@ 5 Denoisers in the Context of Inverse Problems
Consider the following formulation of a linear inverse problem [^16] While we’ll only described linear inverse problems in this exposition, the RED and other frameworks are equally applicable to nonlinear inverse problems (e.g., [82]), albeit with the caveat that some of the nice convexity properties of the overall loss no longer hold.: The data is given by the following model
$\mathbf{y}=H\mathbf{x}+\mathbf{e},$ (63)
where $H\in\mathbb{R}^{m\times n}$ is the forward operator (e.g., degradation or measurements operator), $\mathbf{e}\in\mathbb{R}^{m}$ is additive white Gaussian noise, and the task is retrieving $\mathbf{x}\in\mathbb{R}^{n}$ from $\mathbf{y}\in\mathbb{R}^{m}$.
A nominal solution can be obtained by solving this optimization problem:
$\hat{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{x}}l(\mathbf{y},\mathbf{x})+\lambda\mathbf{R}(\mathbf{x},\alpha),$ (64)
@@
More recently, and independently of the machine learning literature, a fascinating connection has emerged between denoising algorithms and inverse problems. Powerful denoising algorithms, particularly those leveraging deep learning, have been shown to implicitly encode strong priors about natural signals. By incorporating these denoisers into the optimization framework, we can effectively leverage their learned priors to achieve state-of-the-art performance in various inverse problems  [94, 95, 96, 97, 61, 98, 99, 100, 101, 102]. This approach effectively blurs the lines between traditional regularization techniques and modern denoising methods, offering a new paradigm for solving inverse problems.
Learning priors from data has a long history starting in the statistical literature with the concept of “empirical Bayes”(see e.g. [31, 52]). More recently, both implicit and explicit methods have been developed to learn the distribution of images [103, 104, 105, 106, 107, 108]. In particular, the vast recent literature on diffusion models is all about mapping a known distribution (typically a multidimensional Gaussian) to an empirical distribution (learned from collections of images in a desired domain)  [69, 76, 73, 78].
@@
$H^{T}(\mathbf{y}-H\mathbf{x})+\lambda\nabla\mathbf{R}(\mathbf{x},\alpha).$ (65)
A key concern is how to compute $\nabla\mathbf{R}(\mathbf{x},\alpha)$. In this respect, classical choices of $\mathbf{R}(.)$ such as $L_{p}$ norms have been fairly convenient and successful; but also shown to have limited power to model natural images [103].
@@
$\mathbf{R}(\mathbf{x},\alpha)=\frac{1}{2}\mathbf{x}^{T}\;\mathbf{L}(\mathbf{x},\alpha)\;\mathbf{x}.$ (66)
In [61], we developed a natural extension of this idea called Regularization by Denoising (RED), where the regularizer is constructed from a more general denoiser $f(\mathbf{x},\alpha)$:
@@
$\nabla\mathbf{x}^{T}(\mathbf{x}-f(\mathbf{x},\alpha))$ $=$ $2\mathbf{x}-\nabla\left[\mathbf{x}^{T}f(\mathbf{x},\alpha)\right]$ (68)
$=$ $2\mathbf{x}-f(\mathbf{x},\alpha)-\nabla f(\mathbf{x},\alpha)\mathbf{x}$ (69)
$=$ $2\mathbf{x}-2f(\mathbf{x},\alpha),$ (70)
@@
$=$ $\lim_{\epsilon\rightarrow 0}\frac{(1+\epsilon)f(\mathbf{x})-f(\mathbf{x})}{\epsilon}=f(\mathbf{x}).$ (72)
Replacing $\nabla\mathbf{R}_{red}(\mathbf{x},\alpha)=\mathbf{x}-f(\mathbf{x},\alpha),$ for the gradient in (65), we have the following expression for the gradient of the objective:
$H^{T}(\mathbf{y}-H\mathbf{x})+\lambda(\mathbf{x}-f(\mathbf{x},\alpha)).$ (73)
@@
$\mathbf{x}_{k+1}$ $=\mathbf{b}+\mathbf{M}\;f(\mathbf{x}_{k},\alpha),$ (75)
where
$\mathbf{b}=\left[H^{T}H+\lambda I\right]^{-1}H^{T}\mathbf{y},\quad\,\mathbf{M}=\lambda\left[H^{T}H+\lambda I\right]^{-1}.$ (76)
Here, $\mathbf{b}$ is the (fixed) linear pseudo-inverse solution and $\mathbf{M}$ is also a fixed matrix. Procedurally, we start with $\mathbf{x}_{0}=\mathbf{y}$, denoise it, and then a linear operator $\mathbf{M}$ is applied and a bias $\mathbf{b}$ is added - this leads to an updated estimate, and the process is repeated. Note that the structure of this iterative process is not altogether different from a denoising diffusion process  [76] where a denoiser is repeatedly applied. In fact, when $H=I$ we see the structure of a bridge diffusion process [48]:
$\mathbf{x}_{k+1}$ $=\frac{1}{1+\lambda}\mathbf{y}+\frac{\lambda}{1+\lambda}f(\mathbf{x}_{k},\alpha).$ (77)
@@
$\mathbf{R}(\mathbf{x},\alpha)=-\log P(\mathbf{x},\alpha).$ (78)
@@
$P(\mathbf{x},\alpha)\propto\exp{\left[-\mathbf{R}(\mathbf{x},\alpha)\right]}.$ (79)
@@
An alternative approach to solving inverse problems is to leverage pretrained denoisers as priors for generating samples from the posterior distribution [45, 43, 114]. Given measurements $\mathbf{y}$, our goal is to generate samples $\mathbf{x}$ that follow the distribution $P(\mathbf{x}|\mathbf{y})$, where the prior $P(\mathbf{x})$ is implicitly defined by the denoiser.
To achieve this, we can adapt the generative sampling strategy from Equation (54) to sample from the posterior distribution $P(\mathbf{x}|\mathbf{y})$ instead of the prior $P(\mathbf{x})$:
$\frac{d\mathbf{x}_{t}}{dt}$ $=-\frac{1}{2}\frac{d\alpha_{t}}{dt}\nabla\log P(\mathbf{x}_{t}|\mathbf{y},\alpha_{t})$ (81)
$=-\frac{1}{2}\frac{d\alpha_{t}}{dt}\left(\nabla\log P(\mathbf{x}_{t},\alpha_{t})+\nabla\log P(\mathbf{y}|\mathbf{x}_{t},\alpha_{t})\right),$ (82)
starting from $\mathbf{x}_{T}\sim P(\mathbf{x}_{T})$. The second equality is given by Bayes rule.
We recognize the first term as the score function $\nabla\log P(\mathbf{x}_{t},\alpha_{t})$, which can be connected to the MMSE denoiser through Tweedie’s formula (21). The second term in (82) quantifies how well the current sample $\mathbf{x}_{t}$ explains the measurements $\mathbf{y}$, but this is generally intractable to compute.
Diffusion Posterior Sampling framework (DPS): One approach to address this intractability is the Diffusion Posterior Sampling framework [114]. DPS approximates the intractable term with $\log P(\mathbf{y}|\mathbf{x}_{t},\alpha_{t})\simeq\log P(\mathbf{y}|\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}],\alpha_{t})$, based on the assumption that $p(\mathbf{x}_{0}|\mathbf{x}_{t})\simeq\delta(\mathbf{x}_{0}-\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}])$.
Considering a linear measurement model as in Equation (63), this approximation leads to:
$\log P(\mathbf{y}|\mathbf{x}_{t})\simeq-\|H\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]-\mathbf{y}\|^{2}.$ (83)
Substituting this into (82), we obtain:
$\frac{d\mathbf{x}_{t}}{dt}=\frac{1}{2}\frac{d\alpha_{t}}{dt}\left(\frac{\mathbf{x}_{t}-\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]}{\alpha_{t}}+\rho_{t}\nabla_{\mathbf{x}_{t}}\|H\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]-\mathbf{y}||^{2}\right),$ (84)
where $\rho_{t}$ is a hyperparameter balancing the influence of the prior and the measurements. In practice, we utilize a denoiser network $f(\mathbf{x}_{t},\alpha_{t})$ to approximate the conditional expectation $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$.
Growing Importance of Denoising Diffusion Models: Denoising diffusion models are rapidly emerging as a powerful tool for solving inverse problems across various domains. This success often stems from combining the strengths of diffusion models with additional approximations or specialized techniques. A growing body of research explores these approaches ([115, 73, 42, 45, 43, 100, 116, 117, 118, 119, 120, 121, 122]; see [123] for a comprehensive review).